# An explicit form of Ingham’s zero density estimate

### v1

@@ 5.3. Proof of Theorem 2
×$\times\left(\mathfrak{g}_{3}\left(T_{0},\sigma_{5},c_{1},\bm{\mathfrak{a}}\right)^{\frac{1}{2}}+\mathfrak{g}_{4}\left(T_{0},\sigma_{5},c_{2},\bm{\mathfrak{b}}\right)^{\frac{1}{2}}\right).$ (95)
Here, $\mathcal{D}_{4}(x_{0},\sigma)$, $\mathcal{U}^{-}(y_{0})$, $\mathfrak{g}_{3}\left(T_{0},\sigma_{5},c_{1},\bm{\mathfrak{a}}\right)$ and $\mathfrak{g}_{4}\left(T_{0},\sigma_{5},c_{2},\bm{\mathfrak{b}}\right)$ are defined by (59), (73), (5.3.3) and (5.3.3), respectively. It remains to estimate the last two terms on the right-hand side of (84). Beginning with the first such term, note that by writing $z=\sigma+\mathrm{i}T$ in the case of the first integral and $z=\sigma+2\mathrm{i}T$ in the case of the second integral, we have $\sigma\in[1/2,\sigma_{4}]$, implying $-1/2<\sigma+c_{1}<0$, 1/$21/2<\sigma+c_{2}<3/$ 2, $|\sigma+c_{1}+\mathrm{i}(u+kT)|\leq 1/2+kT+\mathfrak{a}_{1}(\log{T})^{\mathfrak{a}_{2}}$ for $k>0$ and $|u|\leq\mathfrak{a}_{1}(\log{T})^{\mathfrak{a}_{2}}$, and $|\sigma+c_{2}+\mathrm{i}(u+kT)|\geq kT-\mathfrak{b}_{1}(\log{T})^{\mathfrak{b}_{2}}$ for $k>0$ and $|u|\leq\mathfrak{b}_{1}(\log{T})^{\mathfrak{b}_{2}}$. Define
@@
Here, $\mathcal{D}_{1}(x_{0},\sigma_{0})$, $\mathcal{U}^{\pm}(y_{0})$, $\mathfrak{g}_{5}\left(T_{0},\sigma_{4},c_{1},\bm{\mathfrak{a}}\right)$ and $\mathfrak{g}_{6}\left(T_{0},\sigma_{4},c_{2},\bm{\mathfrak{b}}\right)$ are defined by (55), (73), (96) and (97), respectively. Turning to the second term on the right-hand side of (84), we have, similarly as before, that $\sigma\in[\sigma_{5},1/2]$, implying $-1<\sigma+c_{1}<0$, $11/2<\sigma+c_{2}<$ 1, $|\sigma+c_{1}+\mathrm{i}(u+kT)|\leq 1+kT+\mathfrak{a}_{1}(\log{T})^{\mathfrak{a}_{2}}$ for $k>0$ and $|u|\leq\mathfrak{a}_{1}(\log{T})^{\mathfrak{a}_{2}}$, and $|\sigma+c_{2}+\mathrm{i}(u+kT)|\geq kT-\mathfrak{b}_{1}(\log{T})^{\mathfrak{b}_{2}}$ for $k>0$ and $|u|\leq\mathfrak{b}_{1}(\log{T})^{\mathfrak{b}_{2}}$. Define
@@
and $\mathcal{J}_{01}\left(T_{0},c_{1},\bm{\mathfrak{a}}\right)$ and $\mathcal{J}_{02}\left(T_{0},c_{2},\bm{\mathfrak{b}}\right)$ are defined by (112) and (114), respectively. In the derivation of (115), we have used inequalities (111) and (113), together with Lemma 19 and $\left|\zeta(1/2+\mathrm{i}t)\right|\leq(2T)^{1/6}\log{(2T)}$ for $t\in[T,2T]$, see [HPY24, Theorem 1.1].

### v2

@@ 5.3. Proof of Theorem 2
$\times\left(\mathfrak{g}_{3}\left(T_{0},\sigma_{5},c_{1},\bm{\mathfrak{a}}\right)^{\frac{1}{2}}+\mathfrak{g}_{4}\left(T_{0},\sigma_{5},c_{2},\bm{\mathfrak{b}}\right)^{\frac{1}{2}}\right).$ (97)
Here, $\mathcal{D}_{4}(x_{0},\sigma)$, $\mathcal{U}^{-}(y_{0})$, $\mathfrak{g}_{3}\left(T_{0},\sigma_{5},c_{1},\bm{\mathfrak{a}}\right)$ and $\mathfrak{g}_{4}\left(T_{0},\sigma_{5},c_{2},\bm{\mathfrak{b}}\right)$ are defined by (61), (75), (5.3.3) and (5.3.3), respectively. It remains to estimate the last two terms on the right-hand side of (86). Beginning with the first such term, note that by writing $z=\sigma+\mathrm{i}T$ in the case of the first integral and $z=\sigma+2\mathrm{i}T$ in the case of the second integral, we have $\sigma\in[1/2,\sigma_{4}]$, implying $1/2<\sigma+c_{1}<0$, $1/2<\sigma+c_{2}<3/2$, $|\sigma+c_{1}+\mathrm{i}(u+kT)|\leq 1/2+kT+\mathfrak{a}_{1}(\log{T})^{\mathfrak{a}_{2}}$ for $k>0$ and $|u|\leq\mathfrak{a}_{1}(\log{T})^{\mathfrak{a}_{2}}$, and $|\sigma+c_{2}+\mathrm{i}(u+kT)|\geq kT-\mathfrak{b}_{1}(\log{T})^{\mathfrak{b}_{2}}$ for $k>0$ and $|u|\leq\mathfrak{b}_{1}(\log{T})^{\mathfrak{b}_{2}}$. Define
@@
Here, $\mathcal{D}_{1}(x_{0},\sigma_{0})$, $\mathcal{U}^{\pm}(y_{0})$, $\mathfrak{g}_{5}\left(T_{0},\sigma_{4},c_{1},\bm{\mathfrak{a}}\right)$ and $\mathfrak{g}_{6}\left(T_{0},\sigma_{4},c_{2},\bm{\mathfrak{b}}\right)$ are defined by (57), (75), (98) and (99), respectively. Turning to the second term on the right-hand side of (86), we have, similarly as before, that $\sigma\in[\sigma_{5},1/2]$, implying $1<\sigma+c_{1}<0$, $1/2<\sigma+c_{2}<1$, $|\sigma+c_{1}+\mathrm{i}(u+kT)|\leq 1+kT+\mathfrak{a}_{1}(\log{T})^{\mathfrak{a}_{2}}$ for $k>0$ and $|u|\leq\mathfrak{a}_{1}(\log{T})^{\mathfrak{a}_{2}}$, and $|\sigma+c_{2}+\mathrm{i}(u+kT)|\geq kT-\mathfrak{b}_{1}(\log{T})^{\mathfrak{b}_{2}}$ for $k>0$ and $|u|\leq\mathfrak{b}_{1}(\log{T})^{\mathfrak{b}_{2}}$. Define
@@
and $\mathcal{J}_{01}\left(T_{0},c_{1},\bm{\mathfrak{a}}\right)$ and $\mathcal{J}_{02}\left(T_{0},c_{2},\bm{\mathfrak{b}}\right)$ are defined by (114) and (116), respectively. In the derivation of (117), we have used inequalities (113) and (115), together with Lemma 19 and $\left|\zeta(1/2+\mathrm{i}t)\right|\leq 0.618t^{\frac{1}{6}}\log t\leq(2T)^{1/6}\log{(2T)}$ for $t\in[T,2T]$, see [HPY24, Theorem 1.1].
@@ 5.3.8. Proof of estimate (13)
6. Appendix
The main purpose of this section is to state an improved zero density estimate from [KLN18]. Lemma 6 can be used instead of [KLN18, Lemma 4.3], which then implies
$N(\sigma,T)\leq\frac{\mathcal{C}_{1}}{2\pi d}T^{\frac{8}{3}\left(1-\sigma\right)}\log(kT)\left(\log{T}\right)^{5-4\sigma}+\frac{\mathcal{C}_{2}}{2\pi d}\log^{2}{T}$
for $T\geq H_{0}=3.0610046\cdot 10^{10}$, where $10^{9}/H_{0}\leq k\leq 1$, $d>0$, and $\mathcal{C}_{1}=\mathcal{C}_{1}\left(\alpha,d,\delta,k,H_{0}-1,\sigma\right)$ and $\mathcal{C}_{2}=\mathcal{C}_{2}\left(d,\eta,k,H_{0}-1,\mu,\sigma\right)$ are functions defined by [KLN18, Equations (4.72) and (4.73)] and depend on several variables, see [KLN18, Lemma 4.14] and [CHJ23, pp. 1223–1224] for details. We choose $k=1$, $\mu=1.2363$ and $\eta=0.25345$. The parameters $\alpha$, $\delta$, and $d$ are calculated such that the value for $\mathcal{C}_{1}\left(\alpha,d,\delta,k,H_{0}-1,\sigma_{2}\right)/(2\pi d)$ is the smallest possible for each of $\sigma_{2}$, see Table 4. The final result is that
$N(\sigma,T)\leq A\cdot T^{\frac{8}{3}\left(1-\sigma\right)}\left(\log{T}\right)^{2(3-2\sigma)}+B\cdot\log^{2}{T}$ (119)
for $T\geq 3\cdot 10^{12}$ and $\sigma\in[\sigma_{1},\sigma_{2}]$, where the values for $A$ and $B$ are given by Table 4.
$\sigma_{1}$ $\sigma_{2}$ AA $B$ $\alpha$ $\delta$ $d$
$0.62500$ $0.65625$ $2.900$ $5.027$ $0.242$ $0.324$ $0.344$
$0.65625$ $0.68750$ $3.474$ $4.789$ $0.224$ $0.324$ $0.343$
$0.68750$ $0.71875$ $4.149$ $4.536$ $0.206$ $0.324$ $0.343$
$0.71875$ $0.75000$ $4.940$ $4.296$ $0.189$ $0.324$ $0.342$
$0.75000$ $0.78125$ $5.860$ $4.042$ $0.171$ $0.324$ $0.342$
$0.78125$ $0.81250$ $6.925$ $3.800$ $0.153$ $0.324$ $0.341$
$0.81250$ $0.84375$ $8.148$ $3.546$ $0.136$ $0.324$ $0.341$
$0.84375$ $0.87500$ $9.543$ $3.301$ $0.119$ $0.324$ $0.340$
$0.87500$ $0.90625$ $11.12$ $3.055$ $0.102$ $0.324$ $0.339$
$0.90625$ $0.93750$ $12.88$ $2.807$ $0.086$ $0.324$ $0.338$
$0.93750$ $0.96875$ $14.81$ $2.558$ $0.070$ $0.325$ $0.337$
$0.96875$ $0.97500$ $15.21$ $2.301$ $0.067$ $0.325$ $0.337$
$0.97500$ $0.98125$ $15.62$ $2.249$ $0.064$ $0.325$ $0.337$
$0.98125$ $0.98750$ $16.04$ $2.204$ $0.061$ $0.325$ $0.336$
$0.98750$ $0.99375$ $16.46$ $2.153$ $0.058$ $0.325$ $0.336$
$0.99375$ $1.00000$ $16.89$ $2.101$ $0.056$ $0.325$ $0.336$
Table 4. The values for $A$ and $B$ from (119) for the corresponding $\sigma\in[\sigma_{1},\sigma_{2}]$.